# On Bits and Bandits: Quantifying the Regret-Information Trade-off

### v2

@@ 3 Regret Lower Bounds Using Fano’s Inequality
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T+1}{\log\mathcal{M}(\rho,\Pi,\epsilon/2)-\log\mathcal{M}(\rho,\Pi,\epsilon)}.$ (5)
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq 1-\frac{\inf_{\delta>0}\left(\log\mathcal{N}_{2}\left(\Pi,\sqrt{\delta}\right)+T\delta\right)+1}{\log\mathcal{M}(\rho,\Pi,\epsilon)}.$ (6)
To obtain regret lower bounds, all we need are the values of the packing and covering numbers of the decision space. Then, we simply need to select a value of $\epsilon$ such that the left-hand side of Equations 4, 5 or 6 is $\frac{1}{2}$ (or any other positive constant). In Appendix C we demonstrate this, by deriving regret lower bounds for several known online problems, which are summarized in Table 1. Moreover, this method can be used to obtain lower bounds for the frequentist regret, due to the mini-max theorem, which states that the worst-case Bayesian regret is equal to the mini-max frequentist regret [Lattimore and Szepesvári, 2019]. More detailed proofs and explanations are provided in Appendix C.
@@ 4 Information Theoretic Bayesian Regret Lower Bounds
Russo and Van Roy [2014, 2016] showed that the regret of Thompson sampling and IDS can be upper bounded by the mutual information.
Theorem 4.1 (Propositions 2,4 of [Russo and Van Roy, 2014] and Propositions 1,5 of [Russo and Van Roy, 2016])).
In the multi-armed bandit with $K$ arms setting, the regret of Thompson sampling and information-directed sampling can be upper bounded by $O(\sqrt{KTI(\pi^{*};\mathcal{H}_{T})})$. Furthermore, if the rewards are a linear function of the decisions, and the decisions are vectors in $\mathbb{R}^{d}$, then the regret is upper bounded by $O\left(\sqrt{dTI(\pi^{*};\mathcal{H}_{T})}\right)$.
Intuitively, in information-theoretic terms, mutual information quantifies the amount of information one random variable contains about another one [Cover, 1999]. This means that mutual information can be utilized to measure the information that is aggregated from various sources using a common measure of bits. Theorem 4.1 shows that we can upper bound the regret when provided with an upper bound for that number of bits. We can now ask the converse question: What is the lower bound of the worst-case Bayesian regret, given $I(\pi^{*};\mathcal{H}_{T})\leq R$? The following proposition answers this question, for problems with a finite decision space.
@@
Let $R\leq H(\pi^{*})$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, such that $I(\pi^{*};\mathcal{H}_{T})\leq R$. Then the worst-case Bayesian regret can be lower bounded by $\Omega\left(\sqrt{TK\frac{R}{\log K}}\right)$. If $\ \Pi\subset\mathbb{B}_{2}^{d}$ and the rewards are a linear function of the decisions, then the worst-case Bayesian regret can be lower bounded by $\Omega\left(\sqrt{Td\frac{R}{\log K}}\right)$.
The proof of Proposition 4.2 utilizes Lemma D.3. This lemma introduces an upper bound for the mutual information, similar to Theorem 3.4, that takes into account the constraint. The rest of the proof follows a similar structure to the proof of Theorem 3.5, substituting the tighter bound into Equation 4. This yields the following Proposition.
@@
Let $R\leq H(\pi^{*})$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, such that $I(\pi^{*};\mathcal{H}_{T})\leq R$. Then,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\gtrsim 1-\frac{2\epsilon^{2}T\frac{R}{\log K}+1}{\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon/2)-\log\mathcal{M}_{1}(\Delta(\Pi),\epsilon)}$ (7)
The rest of the proof follows by selecting the appropriate value of $\epsilon$. Detailed proofs are found in Appendix D. Now, we can quantify the relationship between information and regret using Theorem 4.1 and Proposition 4.2. Proposition 4.2 states that when more bits of information can be accumulated, due to the properties of the environment or the agent, more regret will be suffered. We now consider two agents in a contextual MAB environment, with $K$ arms. The first agent ignores the contextual information and plays each context independently. Thus, for this agent, the mutual information between the optimal policy $\pi^{*}$ and the history $\mathcal{H}_{T}$ is $I_{1}(\pi^{*};\mathcal{H}_{T})\leq\log K$. The second agent utilizes the information in an efficient manner, which reduces the uncertainty in the optimal arm, leading to $I_{2}(\pi^{*};\mathcal{H}_{T})\leq 1$. The amount of contextual information provided is $\log\left(\frac{K}{2}\right)$ bits. From Proposition 4.2 and Theorem 4.1 we know that the optimal behavior of the regret is $\Theta\left(\sqrt{TI(\pi^{*};\mathcal{H}_{T}})\right)$ (omitting terms that depend on the size of the decision space). Thus, the regret of the first agent is $\sqrt{T\log K}$ while for the second one, it is $\sqrt{T}$. Thus, $\log\left(\frac{K}{2}\right)$ bits of external information can be traded with $\sqrt{T}\left(\sqrt{\log K}-1\right)$ regret.
(a) Bayesian regret
(b) Accumulated information
We can also quantify the relationship between information accumulation and regret. Suppose we have two algorithms in the MAB setting, with 16 arms and a prior such that $H(\pi^{*})=4$. During the first 10 episodes, one algorithm accumulates 2 bits of information, while the other accumulates a single bit. In the subsequent 100 episodes, both algorithms accumulate the remaining bits. Since the first algorithm accumulated 2 bits initially, it suffers a regret of $\sqrt{200}$ in these 100 episodes. The second algorithm suffers a regret of $\sqrt{300}$ since it needs to accumulate 3 additional bits. Thus, the second algorithm suffers more regret, even though both accumulate the same number of bits by the end.
@@ 5.1 Stochastic Bayesian bandit
We begin by demonstrating the trade-off in a simple Bayesian MAB problem. In this problem, our model space contains $K$ possible models. The decision space of the problem also contains $K$ decisions, which are called arms. The mean of all arms in every model is $0.5(1-\varepsilon)$, except for one arm for which the mean is $0.5(1+\varepsilon)$. The optimal arm is different for every model. The reward of each arm is Bernoulli distributed. We also fix $K=8,\varepsilon=0.1$. We choose this setting since all of the problems have the same structure. Thus, any difference in regret stems from differences in the prior. In the first experiment, we compare the Bayesian regret obtained by Thompson sampling with different priors for every problem. This experiment allows us to isolate the effect of the total bits on the regret. Results are presented in Figure 2. We see that when fewer bits of information can be accumulated, fewer regret is suffered.
The next experiment is performed under the same setting. Now we assume a uniform prior and compare three different bandit algorithms - EXP3 [Lattimore and Szepesvári, 2020], APS [Xu and Zeevi, 2023], and Thompson sampling [Thompson, 1933]. We also estimate the bits of accumulated information for each algorithm. Results are presented in Figure 3. We can now see how measuring the accumulated bits of information also quantifies the relationship in regret between different bandit algorithms. Additional experimental details are provided in Appendix E.
@@ 5.2 Question Answering with Large Language Models
In the following experiment, we show how the regret-information trade-off can be utilized in a practical setting. We consider a sequential multiple choice question-answering (MCQA) task, where every round we need to provide an answer for a question, out of four possible answers. We also have access to two large language models (LLM), where one has significantly more parameters than the other. Every turn we need to choose which LLM should answer the question. As contextual information, every round we are provided with the question, possible answers, and the output of the small LLM for the given prompt. We receive a positive reward of 1 for answering the question correctly. If we choose to deploy the large LLM we incur a negative reward of 0.1. Our goal is to minimize the accumulated regret.
Since the LLM outputs a probability distribution over tokens, we can measure the information the small LLM will gain after answering the question, in bits. We prompt the small LLM with the question and possible answers, which returns scores for each token. We take the scores corresponding only for the tokens [A, B, C, D], each corresponding to a different answer, and use them to get a distribution over these tokens. The number of information bits is the KL divergence between this distribution and the uniform one. Now we create a bits-based policy, which deploys the large LLM only if the mutual information is below some threshold. This policy queries the large LLM only when a substantial amount of bits can be acquired. We compare the mean regret over a horizon of 200 steps between this policy, and one that randomly selects which LLM to use. The selected threshold value for the bits-based policy also ensures that we query the large LLM the same number of times as the random policy.
@@
Table 2: Mean regret for both policies for a horizon of 200 episodes, under different deployment percentages of the large LLM. The left column denotes the percentage of queries to the large LLM. We report the standard error of the mean.
% Large LLM Bits-based Random
$50\%$ $\mathbf{30.1\pm 0.5}$ $32.5\pm 0.2$
$70\%$ $\mathbf{22.7\pm 0.5}$ $24.8\pm 0.1$
$90\%$ $\mathbf{15.5\pm 0.3}$ $17.1\pm 0.2$
Figure 4: Regret of random and bits-based selection in the MCQA with LLMs task. The random and bits-based policies were allowed to use the large LLM for only 90% of the episodes. Both policies were evaluated across 10 different seeds, with the same questions. The shadowed areas correspond to 2-sigma error bars.
200 questions were randomly selected in advance, which we used to evaluate both agents with 10 different seeds while maintaining the same questions and order. We then repeated this process for 10 different randomly generated sets of questions. We used Mistral 7B Jiang et al. [2023] for the small LLM, and Mixtral 7Bx8 [Jiang et al., 2024] for the large one. For both models, 4-bit quantization [Dettmers and Zettlemoyer, 2023] and flash attention Dao et al. [2022] were applied. Figure 4 compares the random and the bits-based policy for one randomly generated set of questions. Table 2 describes the mean regret of every policy under a different number of large LLM deployments. From the results, we see that deciding whether to query the large LLM or not using bits is better than random selection. Furthermore, this improvement becomes more significant as we increase the number of deployments allowed. This demonstrates how we can easily utilize the quantification of information in bits in online settings. Additional details are provided in Appendix E. Code is available at https://github.com/itaishufaro/bitsandbandits.
@@ 6 Related Work
Bayesian setting:
In the Bayesian setting, prior information is assumed to be expressed as a probability distribution, called the prior. New information is then incorporated using Bayesian inference to update the prior probability, resulting in a posterior distribution. Bayesian algorithms have been extensively studied in multiple online decision-making problems such as multi-armed bandits [Russo and Van Roy, 2014, Kaufmann et al., 2012] and reinforcement learning [Guez et al., 2012, Ghavamzadeh et al., 2016]. Upper bounds for Bayesian algorithms have been proved for both the frequentist [Kaufmann et al., 2011] and Bayesian settings [Russo and Van Roy, 2014, 2016]. Prior dependent regret bounds for Bayesian algorithms have also been studied [Bubeck and Liu, 2013, Russo and Van Roy, 2014]. Previous studies have explored prior-dependent lower bounds for Bayesian regret, but these are constrained to specific forms of priors, such as Gaussian priors [Atsidakou et al., 2024].
Contextual information:
@@
have been employed to derive upper bounds for various online tasks. Russo and Van Roy [2016] introduced an information-theoretic method for establishing Bayesian regret upper bound for Thompson sampling in the multi-armed bandit setting, which depends on the entropy of the prior. This approach was also used to design new Bayesian algorithms that utilize mutual information called information-directed sampling (IDS) [Russo and Van Roy, 2014]. IDS and the information-theoretic method it utilizes were also extended to other tasks such as contextual bandits [Neu et al., 2022], sparse linear bandits [Hao et al., 2021], non-stationary bandits [Min and Russo, 2023, Liu et al., 2023], reinforcement learning [Lu and Van Roy, 2019], non-linear control [Kakade et al., 2020], partial monitoring [Lattimore and Szepesvári, 2019], and other online optimization problems [Liu et al., 2018, Dong et al., 2019, Lattimore and Gyorgy, 2021, Russo and Van Roy, 2018]. The bounds presented in works such as [Russo and Van Roy, 2016, 2014, Neu et al., 2022] are based on the upper bounding of the information ratio. Our work uses a different approach, which obtains nearly matching prior-dependent lower bounds for Thompson sampling and information-directed sampling (IDS). Furthermore, these bounds can be applied beyond the bandit setting. Seldin et al. [2011] used a different information-theoretic method. They utilized PAC-Bayes bounds to obtain information-theoretic upper bounds on the per-round regret, that depend on mutual information.
7 Conclusion and Future Work

### v3

@@ 3 Regret Lower Bounds Using Fano’s Inequality
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2\epsilon^{2}\bar{A}T+1}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)}\right].$ (8)
@@
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{\inf_{\delta>0}\left(\log\mathcal{N}\left(\rho,\Pi,\sqrt{\delta}\right)+T\delta\right)+1}{\log\mathcal{M}(\rho,\Pi,\epsilon)}\right].$ (9)
The proof of Theorem 3.4 is provided in Appendix C.2. To obtain regret lower bounds, all we need are the values of the packing or local packing, and covering numbers of the decision space. Then, we simply need to select a value of $\epsilon$ such that the right-hand side of Equations 7, 8 or 9 is a positive constant. In Appendix C.3 we demonstrate this, by deriving regret lower bounds for several known online problems, which are summarized in Table 1. Moreover, this method can be used to obtain lower bounds for the frequentist regret, due to the mini-max theorem, which states that the worst-case Bayesian regret is equal to the mini-max frequentist regret [Lattimore and Szepesvári, 2019].
@@ 4 Information Theoretic Bayesian Regret Lower Bounds
4.1 Mutual Information Constraint
We now derive Bayesian regret bounds under a constraint on the total amount of information that the agent accumulates. Intuitively, in information-theoretic terms, mutual information quantifies the amount of information one random variable contains about another one [Cover, 1999]. This means that mutual information can be utilized to measure the information that is aggregated from various sources using a common measure of bits. We are interested in lower bounding the regret, under a constraint on the amount of information the agent accumulates. Then, we ask the following question: What is the lower bound of the worst-case Bayesian regret, when the agent accumulates less than $R$ bits of information? We will use the method described in Section 3 to derive regret lower bounds under this constraint. We note that in general $I(\pi^{*};\mathcal{H}_{T})=D\left(P^{\pi^{*}\mid\mathcal{H}_{T}},P^{\pi^{*}}\right)$. Thus, for general agents, $D\left(p_{t},P^{\pi^{*}}\right)$ is the information accumulated up to round $t$, in bits. The following proposition answers this question, for problems with a finite decision space.
@@ 4.1 Mutual Information Constraint
Let $0<R\leq\log K$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, and that the agent accumulates no more than $R$ bits. Then the worst-case Bayesian regret can be lower bounded by $\Omega\left(\sqrt{TK\frac{\log K}{R}}\right)$. Additionally, if $\ \Pi\subset\mathbb{B}_{2}^{d}$ and the rewards are a linear function of the decisions, then the worst-case Bayesian regret can be lower bounded by $\Omega\left(\sqrt{Td\frac{\log K}{R}}\right)$. Furthermore, if $R=0$ the worst-case regret is linear.
The proof of Proposition 4.1 utilizes Lemma D.1, which, similarly to the proof of Theorem 3.3, introduces an upper bound for the mutual information, but this time taking the information constraint into account. The rest of the proof follows a similar structure to the proof of Theorem 3.4, substituting the tighter bound into Equation 7. This yields the following proposition.
@@
Let $R\leq\log K$ be given. Assume that the Bayesian decision problem has a finite decision space of size $K$, such that $D\left(p_{t},P^{\pi^{*}}\right)\leq R$ for all $t$. Then,
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2\epsilon^{2}\bar{A}T\frac{R}{\log K}+1}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Delta(\Pi),\epsilon)}\right]$ (10)
Furthermore, if the decision-making problem is a part of a parametric set, then
$\frac{\mathcal{BR}^{*}(T)}{T\epsilon}\geq\frac{1}{2}\left[1-\frac{2\epsilon^{2}\bar{A}T\frac{R}{\log K}+1}{\log\mathcal{M}^{\mathrm{loc}}(\rho,\Pi,\epsilon)}\right]$ (11)
The proof follows from Proposition 4.2, by selecting an appropriate value of $\epsilon$. We now state an upper bound for MAB problems where the agent accumulates $R$ bits.
Proposition 4.3.
Assume an $K$-MAB problem with no constraints. Let $\tilde{O}_{t},\tilde{\mathcal{H}}_{T}$ be constrained observations and history such that $I(\pi^{*};\tilde{\mathcal{H}_{T}})=R$. The regret of a policy that utilizes Thompson sampling with these observations is upper bounded by $O\left(\log K\sqrt{\frac{KT}{R}}\right)$. Additionally, if the MAB is linear the regret can be upper bounded by $O\left(\log K\sqrt{\frac{dT}{R}}\right)$.
Detailed proofs of Propositions 4.1, 4.2, and 4.3 can be found in Appendix D.
Now, we can quantify the relationship between information and regret. Consider two agents, $A$ and $B$, that play on the same online decision-making task. $A$ has an advantage over $B$, since he already played multiple rounds before. $B$ can query external information to reach the same level of performance as $A$ on the task. Using the regret upper bound in Proposition 4.3 we can get a lower bound on the information $B$ needs. Similarly, Proposition 4.1 gives an upper bound on the number of bits. Thus, Propositions 4.1 and 4.3 quantify information in terms of regret. We note that this quantification is not tight, as the upper bounds differ from the lower bounds by a factor of $\sqrt{\log K}$.
These Propositions also allow us to quantify the relationship between the rate of information accumulation and regret. If both parties play for the same number of rounds, the one that accumulates information faster will also suffer less regret.
4.2 Entropy Constraint
While Proposition 4.1 shows how we can lower bound regret under a constraint on accumulated information, we now shift our focus to a constraint on the prior only, i.e $H(\pi^{*})\leq R$. Russo and Van Roy [2014, 2016] showed that in this scenario, the regret of Thompson sampling and information-directed sampling (IDS) can be upper bounded using $H(\pi^{*})$.
Theorem 4.4 (Propositions 2 and 4 of Russo and Van Roy [2014]).
In the $K$-MAB setting, the regret of Thompson sampling and IDS is upper bounded as $O(\sqrt{KTH(\pi^{*})})$. Furthermore, if the rewards are a linear function of the decisions, and the decisions are vectors in $\mathbb{R}^{d}$, then the regret is upper bounded by $O\left(\sqrt{dTH(\pi^{*})}\right)$.
We next develop an entropy-dependent Bayesian regret lower-bound for algorithms that obey the following assumption.
Assumption 4.1.
$D\left(p_{t},P^{\pi^{*}}\right)\geq c\bar{A}R$ for all $t$, for some constant $c>0$.
Intuitively, this assumption means that we consider only agents that gather a minimal amount of information regarding the problem. We now state the following theorem.
Proposition 4.5.
For any agent for the $K$-MAB problem that satisfy Assumption 4.1, the regret is lower-bounded as $\Omega\left(\sqrt{\frac{KH(\pi^{*})T}{\log K}}\right)$. Additionally, if the MAB is linear, then the regret is lower bounded by $\Omega\left(\sqrt{\frac{dH(\pi^{*})T}{\log K}}\right)$.
The proof of Proposition 4.5 can be found in Appendix D.3. We note that this lower bound behaves like the upper bound Presented in Theorem 4.4, up to a factor of $\sqrt{\log K}$. We also show in the appendix that Assumption 4.1 holds for any algorithm that learns the optimal decision.
@@ 5.1 Stochastic Bayesian Bandit
We begin by demonstrating the trade-off in a simple Bayesian MAB problem. In this problem, our model space contains $K$ possible models. The decision space of the problem also contains $K$ decisions, which are called arms. The mean of all arms in every model is $0.5(1-\varepsilon)$, except for one arm for which the mean is $0.5(1+\varepsilon)$. The optimal arm is different for every model, and the reward of each arm is Bernoulli distributed. We also fix $K=8,\varepsilon=0.1$ and a uniform prior over the problems and compare three different bandit algorithms - EXP3 [Lattimore and Szepesvári, 2020], APS [Xu and Zeevi, 2023], and Thompson sampling [Thompson, 1933]. We also estimate the accumulated information for each algorithm. Results are presented in Figure 2. We see how algorithms that accumulate information quickly also suffer less regret. We also compare Thompson sampling under different priors $H(\pi^{*})$. The results of this experiment can be found in Figure 3. Additional experimental details are provided in Appendix E.
@@ 5.2 Question Answering with Large Language Models
In the following experiment, we show how the regret-information trade-off can be utilized in a practical setting. We consider a sequential multiple choice question-answering (MCQA) task, where every round we need to answer a question, given four possible answers. We also have access to two large language models (LLM), where one has significantly more parameters than the other. At every turn, we choose which LLM should be used to answer the question. Every round we are provided with the question, possible answers, and the output of the small LLM for the given prompt. We receive a positive reward of 1 for answering the question correctly. If we choose to deploy the large LLM we also incur a negative reward of 0.1. Our goal is to minimize the accumulated regret.
On the one hand, the larger LLM is more accurate and will output the correct answer with a higher probability. On the other hand, we do not wish to query the large LLM due to the penalty we suffer, if the small LLM already outputs the correct answer with high probability. We present the following method for selecting the LLM using bits: Use the small LLM to quantify the amount of information that can be obtained. If it is above some threshold, we opt to use the large LLM since it provides more information, which results in less regret as we have seen in Section 3. Otherwise, use the small LLM. We call this approach the bits-based policy.
Since the LLM outputs a probability distribution over tokens, we can measure the information the small LLM will gain after answering the question, in bits. We prompt the small LLM with the question and possible answers, which returns scores for each token. We take the scores corresponding only for the tokens [A, B, C, D], each corresponding to a different answer, and use them to get a distribution over these tokens. We measure the information in bits by calculating the KL divergence between this distribution and the uniform one. We compare the mean regret over a horizon of 200 steps between this policy, and one that randomly selects which LLM to use. The threshold value for the bits-based policy is selected to ensure that we query the large LLM the same number of times as the random policy.
@@
Table 2: Mean regret for both policies for a horizon of 200 episodes, under different deployment percentages of the large LLM. The left column denotes the percentage of queries to the large LLM. For all tables, we used Mixtral-8x7B as the large LLM. For the small LLM, we used Mistral-7B (top table) and Gemma-2B (bottom table). Comparison with other LLMs is provided in Appendix F.
% Deployment Bits-based Random
Mistral 7B $50\%$ $\mathbf{36\pm 1}$ $38.7\pm 0.3$
$70\%$ $\mathbf{26.6\pm 0.8}$ $30.5\pm 0.4$
$90\%$ $\mathbf{21.8\pm 0.4}$ $22.4\pm 0.5$
Gemma 2B $50\%$ $\mathbf{38\pm 2}$ $44.5\pm 0.7$
$70\%$ $\mathbf{32\pm 1}$ $35.6\pm 0.8$
$90\%$ $\mathbf{26\pm 1}$ $27\pm 1$
We used 10 different seeds to generate 10 sets of 200 questions from MMLU randomly. For the small LLM we either used Mistral 7B Jiang et al. [2023], Falcon 7B [Almazrouei et al., 2023], or Gemma 2B [Team et al., 2024]. We used Mixtral 7Bx8 [Jiang et al., 2024], Llama3-70B [AI@Meta, 2024], or Gemma 7B for the large LLM. We applied 4-bit quantization [Dettmers and Zettlemoyer, 2023] for all models and flash attention [Dao et al., 2022] for all models excluding Falcon 7B. Table 2 describes the mean regret of every policy under a different number of large LLM deployments. Tables 4 and 5 provide additional results for different combinations of small and large LLMs.
From the results, we see that deciding whether to query the large LLM or not using bits is better than random selection. Furthermore, this improvement becomes more significant as we increase the number of deployments allowed. This demonstrates how we can easily utilize the quantification of information to improve performance in online tasks. Additional details and results for other models are provided in Appendix F. Code is available here [^1] https://github.com/itaishufaro/bitsandbandits.
@@ 6 Related Work
Bayesian setting.
In the Bayesian setting, prior information is assumed to be expressed as a probability distribution, called the prior. New information is then incorporated using Bayesian inference to update the prior probability, resulting in a posterior distribution. Bayesian algorithms have been extensively studied in multiple online decision-making problems such as multi-armed bandits [Russo and Van Roy, 2014, Kaufmann et al., 2012] and reinforcement learning [Guez et al., 2012, Ghavamzadeh et al., 2016]. Upper bounds for Bayesian algorithms have been proved for both the frequentist [Kaufmann et al., 2011] and Bayesian settings [Russo and Van Roy, 2014, 2016]. Prior dependent regret bounds for Bayesian algorithms have also been studied [Bubeck and Liu, 2013, Russo and Van Roy, 2014]. Previous studies have explored prior-dependent lower bounds for Bayesian regret, but these are limited to specific forms of priors, such as Gaussian priors [Atsidakou et al., 2024].
Contextual information.
@@
have been employed to derive upper bounds for various online tasks. Russo and Van Roy [2016] introduced an information-theoretic method for establishing Bayesian regret upper bound for Thompson sampling in the multi-armed bandit setting, which depends on the entropy of the prior. This approach was also used to design new Bayesian algorithms that utilize mutual information called information-directed sampling (IDS) [Russo and Van Roy, 2014]. IDS and the information-theoretic method it utilizes were also extended to other tasks such as contextual bandits [Neu et al., 2022], sparse linear bandits [Hao et al., 2021], non-stationary bandits [Min and Russo, 2023, Liu et al., 2023], reinforcement learning [Lu and Van Roy, 2019], non-linear control [Kakade et al., 2020], partial monitoring [Lattimore and Szepesvári, 2019], and other online optimization problems [Liu et al., 2018, Dong et al., 2019, Lattimore and Gyorgy, 2021, Russo and Van Roy, 2018]. The bounds presented in works such as [Russo and Van Roy, 2016, 2014, Neu et al., 2022] are based on the upper bounding of the information ratio. Our work uses a different approach, which obtains nearly matching prior-dependent lower bounds for Thompson sampling and information-directed sampling (IDS). Furthermore, these bounds can be applied beyond the bandit setting. Seldin et al. [2011] utilized PAC-Bayes bounds to obtain information-theoretic upper bounds on the per-round regret, that depend on mutual information. Arumugam and Van Roy [2021, 2022] have explored a different information-theoretic method, utilizing rate-distortion theory to minimize regret.
7 Conclusions and Future Work